# Orthogonal projections of hypercubes

### v3

@@ V.1 Analytical investigation
We extend our consideration to a more general distribution by perturbation theory of quantum mechanics. When the distribution is approximately quadripolar, i.e.,
@@
When the two weighted vertices are perpendicular to each other, $\bm{\xi}_{1}^{\top}\bm{\xi}_{2}=0$, the PC2 score linearly depends on the Hamming distance from the second most weighted vertices, similar to the PC1 score. The derivation is in Appendix C.
@@ V.2 Numerical validation
With the ideal distribution of Eq. (11), we show that the PC loading corresponds to the weighted vertices. This argument is expected to apply to distributions approximately similar to Eq. (11), but how valid is this assumption?
@@
where $u_{j;i}$ is the $j$th element of the $i$th PC loading. If all elements of PC loading are the same, the standard deviation is zero.
In Fig. 6, we show the standard deviation dependency on the weight parameter $z$. At $z=\frac{1}{2^{N}}$, where all vertices have the same weight, the standard deviation of PC1 loading abruptly changes to zero, indicating that PC1 is aligned with the selected state $\bm{\xi}$ (or $-\bm{\xi}$). This indicates that as long as the distribution of the vertices is qualitatively similar to Eq. (11), the PC1 is aligned with the selected state $\bm{\xi}$ (or $-\bm{\xi}$). We find that the last PC also shows a sudden change in the standard deviation at $z=\frac{1}{2^{N}}$. The opposite happens to the last PC; the last PC is aligned with the selected state when the selected state is weighted less than the others. Thus, the PC1 is robustly equivalent to the weighted vertices as long as they are weighted more than the others.
Figure 6: The standard deviation of PC loading $\Delta u_{j;i}$ dependency on the weight parameter $z$. We performed PCA on the vertices of the six-dimensional hypercube varying the weight $z$ of the selected state $\bm{\xi}$. The selected vertices are $\bm{\xi}=\begin{bmatrix}+&+&+&+&+&+\end{bmatrix}^{\top}$ and $-\bm{\xi}=\begin{bmatrix}-&-&-&-&-&-\end{bmatrix}^{\top}$. Vertical dashed lines indicate $z=\frac{1}{2^{6}}$, where all vertices have the same weight. For clarity of visualization, we plot a limited number of sampled points of the weight parameter $z$, but we also sampled more densely and confirmed the same trend.
Figure 7: Quality of various projections of hypercubes. The inner product in the original space $\bm{s}^{\top}\bm{s}^{\prime}$ is plotted as a function of the inner product in the projected space $\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}^{\prime}\right)$. (a) The isometric projection of a cube in Fig. 2(a). (b) The isometric projection of a tesseract in Fig. 2(b). (c) The Hamming projection of a cube in Fig. 2(c). (d) The Hamming projection of a tesseract in Fig. 2(d). (e) The fractal projection of a six-dimensional hypercube in Fig. 2(e). (f) The fractal projection of an eight-dimensional hypercube in Fig. 2(f). (g) The projection by PC1 and PC2 of a randomly weighted tesseract in Fig. 4(a). (h) The projection by PC1 and PC2 of a bipolarly weighted tesseract in Fig. 4(f). (i) The projection by PC1 and PC2 of a sexapolarly weighted tesseract in Fig. 5(a). (j) The projection by PC1 and PC3 of a sexapolarly weighted tesseract in Fig. 5(b). The dashed line in each panel crosses the origin with slope one. In (g–j), we show the squared sum of unexplained variance as a shade for the confidence interval. The width of the shade is $2\sum_{i\notin\left\{1,2\right\}}\lambda_{i}$ for (g–i), and $2\sum_{i\notin\left\{1,3\right\}}\lambda_{i}$ for (j). The root mean squared inner-product error is shown at the bottom right. Note that before the calculation of the inner product, we centered all variables to the origin, i.e., $\bm{s}\leftarrow\bm{s}-\left<\bm{s}\right>$ and $\bm{r}\left(\bm{s}\right)\leftarrow\bm{r}\left(\bm{s}\right)-\left<\bm{r}\left(\bm{s}\right)\right>$.
@@ VI.1 Quality of the projections
In Fig. 7, we show the inner products of all possible pairs of vertices in the original space as a function of those in the projected space. In general, the difference between the inner product in the original space and those in the projected space becomes larger as the inner product in the projected space approaches zero, even if the projection method is not PCA. Exceptions are found in the fractal projections [Figs. 2(e) and 2(f)], where the inner products in the original space and projected space do not match even if the inner product in the projected space is relatively larger. All the projections show a similar trend that the inner-product error and the error created by fractal projections is larger than the others. The inner product becomes zero in two conditions: when two vectors are orthogonal, or when one or both of them are zero vectors. As Fig. 5(b) shows, the latter contributes more than the former to the inner-product error because if both vertices of a pair have large probability, the inner-product error becomes large. In fact, the inner-product error increases (cf. inset of Figs. 7(i) and 7(j)) when the weighted vertices are projected around the origin [Figs. 5(b)] compared to when they are projected far from the origin [Figs. 5(a) and 7(i)].
@@ VI.2 Centrality of the projection
In general, the distribution of hypercubic vertices along a linear projection axis can be shown to be Gaussian under fairly weak assumptions. First, note that any transformation vector $\bm{v}$ can be represented as the normalization $\bm{v}=\frac{\bm{w}}{\left|\bm{w}\right|}$ of a weighted superposition of non-overlapping binary sub-states,
$\bm{w}=\sum_{g=1}^{M_{\mathrm{G}}}\frac{a_{g}}{\sqrt{n_{g}}}\bm{\xi}_{g}.$ (25)
Here, $M_{\mathrm{G}}$ is the number of sub-states, reflecting the complexity of the original transformation vector, and $\bm{\xi}_{g}\in\left\{+1,-1,0\right\}^{N}$ is the $g$th sub-state with $n_{g}$ non-zero elements and a weight $a_{g}$. Note that, as $\left|\bm{\xi}_{g}\right|=\sqrt{n_{g}}$ by definition and the sub-states do not overlap with each other, $\bm{\xi}_{g}^{\top}\bm{\xi}_{g^{\prime}}=n_{g}\delta_{g,g^{\prime}}$, the squared norm of $\bm{w}$ is the sum of the squared weights: $\left|\bm{w}\right|^{2}=\sum_{g=1}^{M_{\mathrm{G}}}a_{g}^{2}$. The projected coordinate for a vertex $\bm{s}$ then is expressed as a weighted superposition of sub-coordinates,
$r=\frac{1}{\left|\bm{w}\right|}\bm{w}^{\top}\bm{s}=\frac{1}{\left|\bm{w}\right|}\sum_{g=1}^{M_{\mathrm{G}}}a_{g}r_{g},$ (26)
@@
$r^{2}\right>_{\varrho\left(r\right)}=\frac{1}{\left|\bm{w}\right|^{2}}\sum_{g=1}^{M_{\mathrm{G}}}a_{g}^{2}\left<r_{g}^{2}\right>_{\varrho\left(r_{g}\right)}=1.$ (28)
@@ VI.3 Inner-product error bounds of projections
$\leq$ $2N-\frac{\sqrt{N}}{2}\left[\left|r_{1}\left(\bm{s}\right)+r_{1}\left(\bm{s}^{\prime}\right)\right|+\left|r_{2}\left(\bm{s}\right)+r_{2}\left(\bm{s}^{\prime}\right)\right|\right]$
@@
which is drawn with a white dotted line in Fig. 8(b). The square shape of the possible region is due to the fact that four vertices are chosen to form a square in the Hamming projection as an idealization. We derive Eqs. (45) and (46) in Appendix D. Thus, only a limited region of the $r_{1}\left(\bm{s}\right)$-$r_{2}\left(\bm{s}\right)$ plane in the Hamming projection can be the projected coordinates.

### v4

@@ V.1 Analytical investigation
For the sexapolar distribution, as considered in Fig. 5, the probability distribution can be approximated as
$p\left(\bm{s}\right)\approx\sum_{\mu=1}^{3}\frac{c_{\mu}}{2}\left(\delta_{+\bm{\xi}_{\mu},\bm{s}}+\delta_{-\bm{\xi}_{\mu},\bm{s}}\right),$ (26)
where $c_{1},c_{2},c_{3}\in\left[0,1\right]$ are the weights of the selected states, sorted in descending order, $c_{1}\geq c_{2}\geq c_{3}\geq 0$, and satisfying $\sum_{\bm{s}}p\left(\bm{s}\right)=\sum_{\mu=1}^{3}c_{\mu}=1$. Assuming that the selected states are mutually orthogonal, $\bm{\xi}_{\mu}^{\top}\bm{\xi}_{\nu}=N\delta_{\mu,\nu}$, the mean vector becomes
$\bm{s}\right>=\sum_{\mu=1}^{3}\frac{c_{\mu}}{2}\left(\delta_{+\bm{\xi}_{\mu},\bm{s}}+\delta_{-\bm{\xi}_{\mu},\bm{s}}\right)\bm{s}=\bm{0},$ (27)
and the covariance matrix becomes
$\bm{\varSigma}=\sum_{\mu=1}^{3}\frac{c_{\mu}}{2}\left(\delta_{+\bm{\xi}_{\mu},\bm{s}}+\delta_{-\bm{\xi}_{\mu},\bm{s}}\right)\left(\bm{s}-\bm{0}\right)\left(\bm{s}-\bm{0}\right)^{\top}=\sum_{\mu=1}^{3}c_{\mu}\bm{\xi}_{\mu}\bm{\xi}_{\mu}^{\top}.$ (28)
Then, the eigenvectors and corresponding eigenvalues of the covariance matrix are
$\lambda_{i}=Nc_{i}$ (29)
and
$\bm{u}_{i}=\frac{1}{\sqrt{N}}\bm{\xi}_{i},$ (30)
as already illustrated in Fig. 5.
Next, we consider a generalized bipolar distribution:
$p\left(\bm{s}\right)\approx c_{+}\delta_{+\bm{\xi},\bm{s}}+c_{-}\delta_{-\bm{\xi},\bm{s}},$ (31)
where $c_{+},c_{-}\in[0,1]$ and $c_{+}+c_{-}=1$. For convenience, we define the mean coefficient as $\overline{c}\coloneqq\frac{1}{2}\left(c_{+}+c_{-}\right)$ and the difference as $\Delta c\coloneqq\frac{1}{2}\left(c_{+}-c_{-}\right)$, so that $c_{+}=\overline{c}+\Delta c$ and $c_{-}=\overline{c}-\Delta c$. The mean vector is given by
$\bm{s}\right>=\left(c_{+}-c_{-}\right)\bm{\xi}=\left(2\Delta c\right)\bm{\xi},$ (32)
and the covariance matrix is given by
$\bm{\varSigma}$ $=c_{+}\left[+\bm{\xi}-\left(2\Delta c\right)\bm{\xi}\right]\left[+\bm{\xi}-\left(2\Delta c\right)\bm{\xi}\right]^{\top}$
$\quad+c_{-}\left[-\bm{\xi}-\left(2\Delta c\right)\bm{\xi}\right]\left[-\bm{\xi}-\left(2\Delta c\right)\bm{\xi}\right]^{\top}$
$=\left(\overline{c}+\Delta c\right)\left[\bm{\xi}\bm{\xi}^{\top}-2\left(2\Delta c\right)\bm{\xi}\bm{\xi}^{\top}+\left(2\Delta c\right)^{2}\bm{\xi}\bm{\xi}^{\top}\right]$
$\quad+\left(\overline{c}-\Delta c\right)\left[\bm{\xi}\bm{\xi}^{\top}+2\left(2\Delta c\right)\bm{\xi}\bm{\xi}^{\top}+\left(2\Delta c\right)^{2}\bm{\xi}\bm{\xi}^{\top}\right]$
$=2\overline{c}\left[1+\left(2\Delta c\right)^{2}\right]\bm{\xi}\bm{\xi}^{\top}+2\Delta c\left(-4\Delta c\right)\bm{\xi}\bm{\xi}^{\top}$
$=\left[2\overline{c}+2\overline{c}\left(2\Delta c\right)^{2}-2\left(2\Delta c\right)^{2}\right]\bm{\xi}\bm{\xi}^{\top}$
$=\left[1-\left(2\Delta c\right)^{2}\right]\bm{\xi}\bm{\xi}^{\top}.$ (33)
Note that $\overline{c}=\frac{1}{2}$ from the normalization condition $\sum_{\bm{s}}p\left(\bm{s}\right)=c_{+}+c_{-}=1$. The covariance matrix is proportional to the outer product of the weighted vertex, $\bm{\varSigma}\propto\bm{\xi}\bm{\xi}^{\top}$. Thus, the eigenvector is the same as Eq. (16), but the eigenvalue decreases as $\Delta c$ deviates from zero, $\lambda_{1}=\left[1-\left(2\Delta c\right)^{2}\right]N$. As the weight of the pair deviates from $\frac{1}{2}$, the covariance matrix approaches the zero matrix. If the distribution is idealized as a unipolar one, $\Delta c\to\frac{1}{2}$, the covariance matrix approaches the zero matrix, $\bm{\varSigma}\rightarrow\bm{0}\bm{0}^{\top}$, which results in a zero eigenvalue and arbitrary orthonormal eigenvectors. This result indicates that the difference in weights, $\Delta c$, reduces the contribution of the weighted vertex to the covariance matrix.
Finally, we consider the most general polar distribution:
$p\left(\bm{s}\right)=\sum_{\mu=1}^{M}\left(c_{+,\mu}\delta_{+\bm{\xi}_{\mu},\bm{s}}+c_{-,\mu}\delta_{-\bm{\xi}_{\mu},\bm{s}}\right),$ (34)
[…]
$\bm{s}\right>=\sum_{\mu=1}^{M}\left(c_{+,\mu}-c_{-,\mu}\right)\bm{\xi}_{\mu}=\sum_{\mu=1}^{M}\left(2\Delta c_{\mu}\right)\bm{\xi}_{\mu}.$ (35)
After some algebra101010 See the Supplemental Material [16] for the detailed derivation. similar to derive Eq. (33), the covariance matrix becomes
$\bm{\varSigma}$ $=\sum_{\mu=1}^{M}\left[2\overline{c}_{\mu}-\left(2\Delta c_{\mu}\right)^{2}\right]\bm{\xi}_{\mu}\bm{\xi}_{\mu}^{\top}$
$\quad-\sum_{\mu=1}^{M}\sum_{\nu=\mu+1}^{M}\left(2\Delta c_{\mu}\right)\left(2\Delta c_{\nu}\right)\left(\bm{\xi}_{\mu}\bm{\xi}_{\nu}^{\top}+\bm{\xi}_{\nu}\bm{\xi}_{\mu}^{\top}\right).$ (36)
If the coefficient differences are zero, $\Delta c_{\mu}=0$ for all $\mu$, the covariance matrix simplifies to
$\bm{\varSigma}=\sum_{\mu=1}^{M}\left(2\overline{c}_{\mu}\right)\bm{\xi}_{\mu}\bm{\xi}_{\mu}^{\top},$ (37)
and, assuming that the selected states are mutually orthogonal, $\bm{\xi}_{\mu}^{\top}\bm{\xi}_{\nu}=N\delta_{\mu,\nu}$, the eigenvectors and eigenvalues are the same as in Eq. (28).
To understand the qualitative behavior, let us consider $M=2$ with $\Delta c_{1}=0$ and $\Delta c_{2}\neq 0$. In this case, the covariance matrix becomes
$\bm{\varSigma}=2\overline{c}_{1}\bm{\xi}_{1}\bm{\xi}_{1}^{\top}+\left[2\overline{c}_{2}-\left(2\Delta c_{2}\right)^{2}\right]\bm{\xi}_{2}\bm{\xi}_{2}^{\top}.$ (38)
If $\bm{\xi}_{1}^{\top}\bm{\xi}_{2}=0$, the eigenvalues are $\lambda\in\left\{2\overline{c}_{1}N,\,\left[2\overline{c}_{2}-\left(2\Delta c_{2}\right)^{2}\right]N\right\}$ and the corresponding eigenvectors are $\bm{u}\in\left\{\frac{1}{\sqrt{N}}\bm{\xi}_{1},\,\frac{1}{\sqrt{N}}\bm{\xi}_{2}\right\}$. Note that even if $\overline{c}_{1}<\overline{c}_{2}$, depending on the value of $\Delta c_{2}$, the eigenvalue corresponding to $\bm{\xi}_{2}$ can be larger than the eigenvalue corresponding to $\bm{\xi}_{1}$. When $\Delta c_{1}\neq 0$ and $\Delta c_{2}\neq 0$, the covariance matrix becomes
$\bm{\varSigma}$ $=\left[2\overline{c}_{1}-\left(2\Delta c_{1}\right)^{2}\right]\bm{\xi}_{1}\bm{\xi}_{1}^{\top}+\left[2\overline{c}_{2}-\left(2\Delta c_{2}\right)^{2}\right]\bm{\xi}_{2}\bm{\xi}_{2}^{\top}$
$\quad-\left(2\Delta c_{1}\right)\left(2\Delta c_{2}\right)\left(\bm{\xi}_{1}\bm{\xi}_{2}^{\top}+\bm{\xi}_{2}\bm{\xi}_{1}^{\top}\right),$ (39)
which makes it challenging to analytically find the eigenvectors and eigenvalues due to the presence of the symmetric outer product matrix $\bm{\xi}_{1}\bm{\xi}_{2}^{\top}+\bm{\xi}_{2}\bm{\xi}_{1}^{\top}$. If $\bm{\xi}_{1}$ and $\bm{\xi}_{2}$ are orthogonal, the eigenvector of this symmetric outer product matrix is $\frac{1}{\sqrt{2N}}\left(\bm{\xi}_{1}+\bm{\xi}_{2}\right)$ with eigenvalue $N$. Thus, the eigenvector of the symmetric outer product matrix is linearly dependent on the weighted vertices, and an analytical derivation of the eigenvectors and eigenvalues remains challenging. This difficulty arises from the nonzero centering by the mean vector.
Our consideration of the generalized bipolar distribution shows that centering with a nonzero mean vector can complicate analytical understanding. It is, however, important to note that the center of the hypercube is a meaningful point of reference for our visualization purposes—if PCA rotates to project the hypercubic vertices around a point other than the center, the resulting leading PC loading becomes biased toward the center of the hypercube. This can be misleading for visualization and should be avoided. We expect that centering at the zero vector is the most appropriate choice for PCA. See also Refs. [93, 114] for a review of centering in PCA.
@@
We further extend our analysis to more general distributions with non-orthonormal weighted vertices, i.e., $\bm{\xi}_{\mu}^{\top}\bm{\xi}_{\nu}\neq 0$, using perturbation theory from quantum mechanics. When the distribution is approximately quadripolar, that is,
@@
When the two weighted vertices are perpendicular to each other, $\bm{\xi}_{1}^{\top}\bm{\xi}_{2}=0$, the PC2 score depends linearly on the Hamming distance from the second most weighted vertex, similar to the PC1 score. The derivation is given in Appendix C.
Figure 6: The standard deviation of PC loading, $\Delta u_{j;i}$, as a function of the weight parameter $z$. We performed PCA on the vertices of a six-dimensional hypercube, varying the weight $z$ assigned to the selected state $\bm{\xi}$. The selected vertices are $\bm{\xi}=\begin{bmatrix}+&+&+&+&+&+\end{bmatrix}^{\top}$ and $-\bm{\xi}=\begin{bmatrix}-&-&-&-&-&-\end{bmatrix}^{\top}$. For each value of $z$, we sampled 1024 realizations of random weights $\left\{c_{\bm{\xi}^{\prime}}\left(z\right)\right\}$ from the uniform distribution, and the shaded area indicates the sample standard deviation. Vertical dashed lines indicate $z=1$, where the selected vertices lose their distinguished weight.
@@ V.2 Numerical validation
With the ideal bipolar distribution of Eq. (13), we have shown that the PC loading corresponds to the weighted vertices. This argument is expected to hold for distributions that are approximately similar to Eq. (13). Indeed, in the previous Sec. V.1, a bipolar distribution with background [Eq. (18)] was introduced, and it was shown that the PC1 loading aligns with the weighted vertices as long as the distribution retains a bipolar-like structure. Nevertheless, how robust is the assumption of a bipolar distribution with background?
@@
where $u_{j;i}$ denotes the $j$th element of the $i$th PC loading. If all elements of a PC loading are identical, the standard deviation is zero.
In Fig. 6, we show the dependence of the standard deviation on the weight parameter $z$. At $z\approx 1$, where some randomly weighted vertices start to overwhelm the of weighted vertices, the standard deviation of the PC1 loading exhibits a nonlinear increase, indicating that PC1 is no longer aligned with the selected state $+\bm{\xi}$ (or $-\bm{\xi}$). The qualitative equivalence of the covariance matrix [Eq. (19)] for the bipolar distribution with background [Eq. (18)] to the covariance matrix of the bipolar distribution [Eq. (15)] holds for the uniform random weight distribution as long as the maximum random weight $z$ is approximately less than that of the selected vertices.
@@ VI Dependency of inner-product error on projected coordinates and its inevitability
Figure 7: Quality of various projections of hypercubes. The inner product in the original space $\bm{s}^{\top}\bm{s}^{\prime}$ is plotted as a function of the inner product in the projected space $\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}^{\prime}\right)$. (a) The isometric projection of a cube in Fig. 2(a). (b) The isometric projection of a tesseract in Fig. 2(b). (c) The Hamming projection of a cube in Fig. 2(c). (d) The Hamming projection of a tesseract in Fig. 2(d). (e) The fractal projection of a six-dimensional hypercube in Fig. 2(e). (f) The fractal projection of an eight-dimensional hypercube in Fig. 2(f). (g) The projection by PC1 and PC2 of a randomly weighted tesseract in Fig. 4(a). (h) The projection by PC1 and PC2 of a bipolarly weighted tesseract in Fig. 4(f). (i) The projection by PC1 and PC2 of a sexapolarly weighted tesseract in Fig. 5(a). (j) The projection by PC1 and PC3 of a sexapolarly weighted tesseract in Fig. 5(b). The dashed line in each panel crosses the origin with slope one. In (g)–(j), we show the squared sum of unexplained variance as a shade for the confidence interval. The width of the shade is $2\sum_{i\notin\left\{1,2\right\}}\lambda_{i}$ for (g)–(i), and $2\sum_{i\notin\left\{1,3\right\}}\lambda_{i}$ for (j). The root mean squared inner-product error is shown at the bottom right. Note that before the calculation of the inner product, we centered all variables to the origin, i.e., $\bm{s}\leftarrow\bm{s}-\left<\bm{s}\right>$ and $\bm{r}\left(\bm{s}\right)\leftarrow\bm{r}\left(\bm{s}\right)-\left<\bm{r}\left(\bm{s}\right)\right>$.
@@ VI.1 Quality of the projections
In Fig. 7, we show the inner products of all possible pairs of vertices in the original space as a function of those in the projected space. In general, the difference between the inner product in the original space and those in the projected space becomes larger as the inner product in the projected space approaches zero, even if the projection method is not PCA. Exceptions are found in the fractal projections [Figs. 2(e) and 2(f)], where the inner products in the original space and projected space do not match even if the inner product in the projected space is relatively larger. All the projections show a similar trend that the inner-product error and the error created by fractal projections is larger than the others. The inner product becomes zero in two conditions: when two vectors are orthogonal, or when one or both of them are zero vectors. As Fig. 5(b) shows, the latter contributes more than the former to the inner-product error because if both vertices of a pair have large probability, the inner-product error becomes large. In fact, the inner-product error increases [cf. inset of Figs. 7(i) and 7(j)] when the weighted vertices are projected around the origin [Fig. 5(b)] compared to when they are projected far from the origin [Figs. 5(a) and 7(i)].
@@ VI.2 Centrality of the projection
In general, the distribution of hypercubic vertices along a linear projection axis can be shown to be Gaussian under fairly weak assumptions. First, note that any transformation vector $\bm{v}$ can be represented as the normalization $\bm{v}=\frac{\bm{q}}{\left|\bm{q}\right|}$ of a weighted superposition of non-overlapping binary sub-states,
$\bm{q}=\sum_{g=1}^{M_{\mathrm{s}}}\frac{a_{g}}{\sqrt{n_{g}}}\bm{\xi}_{g}.$ (46)
Here, $M_{\mathrm{s}}$ is the number of sub-states, reflecting the complexity of the original transformation vector, and $\bm{\xi}_{g}\in\left\{+1,-1,0\right\}^{N}$ is the $g$th sub-state with $n_{g}$ non-zero elements and a weight $a_{g}$. Note that, as $\left|\bm{\xi}_{g}\right|=\sqrt{n_{g}}$ by definition and the sub-states do not overlap with each other, $\bm{\xi}_{g}^{\top}\bm{\xi}_{g^{\prime}}=n_{g}\delta_{g,g^{\prime}}$, the squared norm of $\bm{q}$ is the sum of the squared weights: $\left|\bm{q}\right|^{2}=\sum_{g=1}^{M_{\mathrm{s}}}a_{g}^{2}$. The projected coordinate for a vertex $\bm{s}$ then is expressed as a weighted superposition of sub-coordinates,
$r=\frac{1}{\left|\bm{q}\right|}\bm{q}^{\top}\bm{s}=\frac{1}{\left|\bm{q}\right|}\sum_{g=1}^{M_{\mathrm{s}}}a_{g}r_{g},$ (47)
@@
$r^{2}\right>_{\varrho\left(r\right)}=\frac{1}{\left|\bm{q}\right|^{2}}\sum_{g=1}^{M_{\mathrm{s}}}a_{g}^{2}\left<r_{g}^{2}\right>_{\varrho\left(r_{g}\right)}=1.$ (49)
@@ VI.3 Inner-product error bounds of projections
$\leq 2N-\frac{\sqrt{N}}{2}\left[\left|r_{1}\left(\bm{s}\right)+r_{1}\left(\bm{s}^{\prime}\right)\right|+\left|r_{2}\left(\bm{s}\right)+r_{2}\left(\bm{s}^{\prime}\right)\right|\right]$
@@
which is drawn with a white dotted line in Fig. 8(b). The square shape of the possible region is due to the fact that four vertices are chosen to form a square in the Hamming projection as an idealization. We derive Eqs. (66) and (67) in Appendix D. Thus, only a limited region of the $r_{1}\left(\bm{s}\right)$$r_{2}\left(\bm{s}\right)$ plane in the Hamming projection can be the projected coordinates.